Emily bought two blouses and a pair of pants. The total amount she paid for the items is not more than 980. What are the prices of each item


1
Expert's answer
2021-02-25T05:05:32-0500

Step-by-step explanation:

2x+y=980

Let x, be blouse = Php 150.00

Let y, for a pair of pants

2x + y = 980

2( Php 150.00) + y = Php 980

300 + y = 980

Y= 980 - 300

Y= Php 680.00 is the cost of a pair of pants